The Communication of Grammatical Proficiency*

M. Trevor Shanklin

0 Introduction

In a recent article, James reviews the function of applied linguistics and concludes that it should serve as an interface between “linguistics and teaching methodology” (1993 : 30). My contention in this paper will be that recent work in generative grammar, on language processing and on error analysis provides a basis for making principled decisions in the classroom concerning the types of skills we wish to impart and the method of imparting them. The applied linguist as someone who is competent in both linguistics and teaching methodology assists in the application of linguistic theory to the classroom and of observations concerning the language learning process to linguistic theory. This latter function becomes all the more important when we realize that grammatical theory itself is simply a tool for understanding the language. Thus this feedback process is a critical link in the practical application of theory to classroom issues concerning language acquisition. In this regard, linguists should aspire to learn more about what is going on in the classroom and teachers to learn more about what is going on in linguistics.^{1}

Combining the three strands of investigation and the two roles alluded to above, we discover (i) that grammatical proficiency is both an important pedagogical skill and an important part of target language proficiency; (ii) that findings within the framework of generative grammar allow us to develop our language awareness to assist in achieving these functions of grammatical proficiency.

1 Definition

148 M. Trevor Shanklin

Grammar is an abstraction, a hypothesis about the way that language works. The theory of generative grammar pays tribute to this exploratory role of grammar and perhaps helps to correct a perception that ‘grammar’ is a stored set of rules that we set out to learn. One exemplification of this is the distinction made between internal (I) and external (E) language, for example in Chomsky (1986). As put by Burton-Roberts (1993), whatever a string of words on paper is, it is not a sentence, i.e. it has “no linguistic properties at all.” What makes that string of words a sentence, are the mental properties projected onto them by the language speaker, the hierarchical arrangement of constituents: “I-language has no life apart from mental representation,” i.e. sentences belong to the I-language and not the E-language. In exploring the grammar, we are exploring those mental properties. This has consequences for second-language teaching: while reference grammars catalog external language data, a generative grammar will suggest means to evaluate issues of appropriacy and acceptability, i.e. the implicit knowledge of the language.

In this paper the term grammatical proficiency will be used to mean the explicit awareness of how language works. The expression ‘proficiency’ is chosen over ‘competence’ to avoid the confusion caused by the competence/performance distinction. Taylor defines proficiency as the “ability to make use of competence.” This is not quite performance which is when “proficiency is put to use” (1988 : 166). Nonetheless the term can thus mean skills that are associated with performance. Proficiency is defined by Taylor as a ‘dynamic concept’, which is assuredly the type of concept we are dealing with in second-language acquisition. At the same time, the understanding seems to fall in with Canale’s distinction between knowledge and skill with regard to communicative competence (1983 : 6). In discussing the applications of his model for second language teaching, Canale further distinguishes ‘knowledge-’ and ‘skill-oriented teaching’ (op.cit. : 14). The type of grammatical competence that we are concerned with in the classroom, in other words, is the skill-oriented competence, which is here given the name grammatical proficiency. The distinction by Canale is a natural consequence of trying to extend our understanding of the language system furthered by research within the generative paradigm to second-language acquisition which is a dynamic process. Taylor stresses that the term competence as used by Chomsky is not an ability but a type of knowledge. In the second-language classroom, teachers are concerned with imparting an ability, a skill, and their effectiveness will in part be determined by the understanding gleaned from the study of language competence, the implicit knowledge of a language.

One disadvantage of the failure to take development into consideration when evaluating competence is the over-reliance in linguistic theory on native-speaker competence. I think non-native speakers can rely much more on their intuitions than they (or we) are aware of.

Communicating Grammatical Proficiency 149

That takes us some way towards understanding the term ‘proficiency’, but what about the term ‘grammatical’? As suggested above, grammatical proficiency will be defined here as the ability to make judgements about the appropriateness/acceptability of an utterance. Such a definition encompasses a large area, including register difference and other variation that belongs to pragmatics or sociolinguistics (and therefore would incorporate grammatical, discourse and sociolinguistic competence in Canale’s scheme). Nonetheless, the definition as it stands is not restrictive enough. While it includes being able to make decisions about the acceptability of a segment of discourse, it does not distinguish this from the introspection of a native-speaker. Yet, if a native speaker can say a sentence is acceptable, without being able to revert to basic grammatical or linguistic notions such as subject and object, we would not consider this grammatically proficient (though certainly competent). The competence of a native speaker that entails an implicit knowledge of the grammar does not yet encompass the explicit awareness that grammatical proficiency suggests. The definition needs to be modified so that explicit awareness is subsumed but not implicit. Let’s say then that:

(1) Grammatical proficiency is the ability to make judgements about the acceptability and appropriateness of an utterance with specific reference to grammatical notions.

In the section below on language processing we will see that this understanding of grammar is similar to Bialystok’s ‘symbolic’ knowledge of language. Left out of the definition as it stands is the mention of rules. There is a vast array of resource material that is in some ways in conflict with each other. Also in conflict are the approaches taken to rules by native and non-native speakers. Whereas native speakers use rules to bolster intuition, non-natives often use rules more religiously as a raft to cling to in the stormy sea of uncertain judgements (cf. Medgyes 1994). As mentioned above, it is my belief that non-natives can rely much more on their own implicit grammar and hence also on introspection. Yet, when disputes arise, the desire to be able to turn to an authoritative work is natural. Here, however, we have to recognize that rule books contradict themselves, native speakers (and linguists!) disagree about the acceptability of sentences/utterances and finally that the broad domain of language will always give rise to questions that have not been codified into rules while the variability of language both synchronically and diachronically leads to continuous modification of/amendment to the rules that we do have.^{2} 1.1 Grammar as Language Awareness

Reference grammars can only be one source of input in reaching decisions about language form. The conflict between rule-books can be illustrated with a brief example. My advanced first year students at CETT in Budapest asked whether in a construction like I’d rather the contracted verb could be interpreted as had as well as would. I thought it could but agreed to look into the matter. First I consulted the Oxford English Dictionary (OED), which makes clear that had rather has had a long and glorious history in the English language. A typical entry there is the following from Shelley:

I had rather err with Plato than be right with Horace

Swan (1980) comments on current usage saying:

“Many grammars and dictionaries say the expression had rather is used in the same way as would rather. This is not true: had rather does not exist is standard Modern English.” (¶518c)

Now that is an authoritative stance. Quirk et al. (1985) take a milder view:

“Sometimes (particularly in AmE) the uncontracted form of ’d rather is realized as had rather […] These variations, and similar variations with would sooner, etc. have presumably arisen because of the ambiguity of the contraction ’d.” (¶3.45 Note b)

I wonder. The historical base of had rather documented in the OED suggests an implicit relationship in the mental grammar between the auxiliary had and rather of a deeper nature than one formed solely via analogy with other would contractions. In terms of judging what is acceptable in standard Modern English, we need more evidence than Swan’s assertion that the construction does not exist. We can invite native speakers to introspect (and I come up with a plausible collocation) or we might consult extensive data banks.

Grammars have to choose between their task as a descriptive pedagogic device and being true to the facts. Once Burton-Roberts, author of the successful Analysing Sentences in responding to the observation that the use of the verbal group does not mesh with the deeper knowledge that we have of English structure, acknowledged that tests for constituent structure would verify this. Nonetheless, it was easier to communicate basic grammatical knowledge, he felt, using the pedagogic device of a ‘verbal group.’ He went on to say that in a higher-level seminar he would then confront students with data that make them revise those earlier assumptions.^{3}

The examples above illustrate the conflict between rule-books but also imply a conflict between rule-books and judgements via introspection.

Another example also illustrates this latter point. When I asked two British colleagues of mine at CETT to choose between the following two sentences, both said they preferred the first:

(2) a. He doesn’t dare do that.

b. He doesn’t dare to do that.

The second was acceptable for them but they would use the first, with the bare infinitive.

Greenbaum & Quirk (1990) mention the difference between the main verb (do-support/to-infinitive) and modal usage (absence of do-support/bare infinitive) of dare, adding the following note: “Blends of the two constructions are widely acceptable: they do not dare ask for me.” (¶3.17. Note)

The Student’s Grammar stops there, but the Comprehensive Grammar goes on to say: “one would expect these to be ungrammatical but they are not.” Then the idiomatic construction is cited:

(3) Don’t you dare speak to me like that.

implying that the to-infinitive here would be inappropriate. Meanwhile, Swan does not even mention this construction (main verb dare and bare infinitive). His only comment is that “in practice dare is not a very common word in Modern English.”^{4}

Thomson & Martinet use the term ‘past subjunctive’ for the use of the past with hypothetical meaning, while Quirk et al. distinguish the hypothetical past from the past subjunctive. The phrasing adopted by Quirk et al. allows us to distinguish a formal concept from a notional one. Thus, for them were is the only past subjunctive, as it is the only verb that shows a difference in form in certain ‘hypothetical’ environments. This discrimination is advantageous, indeed necessary, as it allows us to keep apart was and were when they appear in similar environments. Calling this use of was the hypothetical past makes good sense, providing a notional motivation for its use. Meanwhile, the subjunctive form were harks back to the wide-spread use of a distinct past subjunctive form. At one point, Thomson & Martinet say that were is used “in expression of doubt or unreality.” A student could draw the conclusion from this that were could be used the way the past subjunctive used in reported speech in Modern German, producing a sentence like

(4) He said he were here.

with the meaning of ‘he said he was here but I don’t believe him’. It is not, in other words, used invariably in “expression of doubt or uncertainty.”

Yet, in general a less precise conflation of terms is preferred to more precise formulations if the material is made more accessible, especially if the audience is less advanced. Trainees say they teach incomplete rules so as not to confuse lower-level learners, knowing that in some sense the rule they are providing is false and that at some higher level the ‘exceptions’ to the earlier formulation will be taught. In fact, key items on the university entrance exam are often the kind that have the student demonstrate his/her knowledge that a construction that is deemed inaccurate at a lower level is in fact accurate given the right linguistic environment or context (such as the use of the conditional in a conditional clause).

If the language learner were to become truly dependent on a rule book, wouldn’t s/he be in a muddle! I think we can draw the following conclusions from the preceding discussion:

(i) Teacher-trainees agree that pupils cannot tolerate cases of ambiguity. On the one hand, we need to honor this as teachers. On the other hand, prescriptive and proscriptive grammars do not seem to be the solution, as they create false assumptions about the body of data (language) being analysed.

(ii) The simplification of grammars to enhance accessibility can lead to

inaccuracies. (iii) Pedagogical grammars are helpful, as are dictionaries and other ref-

erences, as a source of data. The real grammar, as made evident by advances in generative theory of the last three decades, is in the mind and it is this grammar that we need to continue to explore.

(iv) In order for this exploration to take place, two pedagogical requirements have to be fulfilled: (a) lateral communication has to take place, here as in other spheres of society; this implies, among other things, relinquishing a dependence on a top-down hierarchical model; (b) learner autonomy needs to be the focal point of the curriculum so that introspection, lateral communication and a judicious selection between ‘authorities’ takes place.

1.2 Generative Theory

Generative grammar has offered the means of relating grammar to language acquisition because the grammar is understood as a subset of psychology, rather than a social science, de Saussure’s position, or as a rule book. After Chomsky first presented this position with the publication of Syntactic Structures in 1956, generative grammar soon became associated with the fledgling discipline of cognitive science. Generative grammar itself has undergone an extensive evolution since then. One of the most important developments has been the articulation of X-bar syntax, a coherent set of rules for drawing syntactic representations. Another important feature that characterizes the Principles and Parameters framework (Chomsky 1982, 1986) is the emphasis on explanatory adequacy as a criterion for subjecting these syntactic representations to. Does, in other words, the representation itself provide an answer to the types of phenomena we observe in language behavior? Given the psychological reality that is associated with a representation in this framework, actual evidence from the second-language classroom should help shed light on the accuracy of the representation at the same time that the representation can help make sense of the grammar exhibited in the classroom.

1.3 Interlanguage

A key concept in the study of second-language acquisition is that of interlanguage and the inherent assumption in the term that every stage of the language learner’s development can be characterized by a grammar that provides the rules for the current system. This in turn shares certain assumptions with the generative framework, that learning a language can be depicted as the postulation of hypotheses, largely unconscious, about how the language system operates. The grammar of the interlanguage reveals the current hypotheses held by the language learner. One means of gaining glimpses into the interlanguage is by looking at the types of errors that language learners make. These errors can be viewed as stemming from first-language interference or as being developmental, independent of the first language. Much work in the 70s was carried out to try to understand second-language acquisition in terms of interference. It seems that the revealing errors in terms of both linguistic theory and understanding the second language acquisition process are developmental rather than those of interference. Using these, it is possible to make the implicit interlanguage grammar explicit.

2 Learning Strategies and Fluency

A variety of research projects in the late 70s and early 80s tried to identify the strategies that good language learners employed to see if they correlated in general with successful language learning. A hypothesis Chaudron cites to explain the results of a number of research projects under review is that a “learner’s involvement in interactive negotiation of meaning” might be a predictor of successful language learning (1988 : 117). This would clearly have an impact on what teachers do in the language classroom. As Bialystok stresses in her review of research into learning strategies, these studies are largely taxonomic and thus fail to make sense out of the strategies studied in terms of a theory of language processing. This failure is seen in the variety of classification used by different authors.

From the standpoint of generative grammar, a complicating factor in discussing language processing is the claim that there is a separate language faculty in the mind where this takes place and therefore must be viewed separately from other cognitive properties. This does not mesh with the position taken by Bialystok. A student at ELTE has proposed a model that posits a surface level where factors like motivation play a role, a second level of specific language properties and a deeper level of general cognitive properties (cf. Turóczi 1994). Perhaps such an approach would make both positions compatible.

Trainees I have worked with often raise the issue of student intelligence or talent for learning languages. My response is to have them think about what those pupils of theirs are doing right and how to impart this skill to others, that is to analyse their learning strategies. Yet the strategies themselves only first make sense in a coherent theoretical framework of what is involved in language learning or language acquisition. As Bygate observes:

“[…] the pedagogical criteria for using a type of activity need to be complemented by an awareness of its psycholinguistic implications […] And even if motivation and self-pacing are the principal criteria for the selection of activities for use in the classroom, then this leaves an enormous choice of activity open to the teacher. Professional judgement therefore needs also to be exercised with regard to language processing skills which particular activities are likely to require and develop, and to the effects these exercises are likely to have on the students’ language learning.” (Bygate 1988 : 79)

A good starting point for understanding language proficiency is the notion of fluency, which, as James points out, teachers often make reference to. Whether grammatical proficiency, understood here as explicit awareness of the language system, constitutes part of fluency is open to debate. At the one end of the spectrum there is either Krashen or the strong Krashen position that “there is no interface between conscious learning and subconscious acquisition” (Schmidt (1992) citing Krashen (1985)). At the other end is Bialystok:

“For Bialystok, the direction of development [of fluency] is from low to high levels of control. Not only the selective allocation of attention but also more specific control processes such as rehearsal, search, planning, monitoring, and decision making of all kinds.” (my emphasis) (Schmidt 1992 : 362)

Schmidt himself lies between these two positions. He indeed says near the beginning of his article on fluency that he is adopting Brumfit’s definition that makes a distinction between fluency and accuracy. For him fluency is above all a “performance phenomenon with particular emphasis on its temporal aspects.” He even agrees with the position taken by Schumann (1990) that the speaker of a ‘pidginized interlanguage […] agrammatic with respect to the target language’ is fluent, and contrasts this understanding with a ‘more global proficiency definition’ of fluency. However, a short while later, Schmidt mentions that in one study a speaker was not judged to be fluent because of accuracy errors and that:

“hearer-based impressions of fluency are holistic, influenced by considerations of accuracy as well as by the temporal, performance aspects on which I will focus in this paper.” (1988 : 359)

But this position is not really adopted in the paper. In his review of work by Bialystok, Schmidt points out that:

“In addition, at least a partial shift in focus away from automaticity toward efficient self-regulation as an essential characteristic of fluency is justified by the fact that skilled performance requires a balance between the speed of automatic processing and the goaldirectedness of controlled processing.” (1988 : 366)

Indeed, citing other studies, Schmidt acknowledges that “a great deal of empirical data suggest that automatic processes […] are subject to attentional control to some degree” (ibidem).

In this state-of-the-art paper, Schmidt reviews a number of models for fluency and it is not always clear what his own position is. At the least, he is interested in investigating the development of fluency via the psychological learning mechanisms “underlying the contrast between automatic and controlled processing” (abstract). This notion of controlled processing is an important addition, in my opinion, to the concept of automatic processing with which fluency has often been equated. Through the review of Bialystok’s and other models, Schmidt shows that practice should be able to influence fluency by making processes automatic and freeing up attention for higher level tasks. For example:

“The nonfluent learners’ pauses, false starts and other signs of hesitation reflect the need to focus attention on the lower levels of planning, whereas fluent learners act more like native speakers in exhibiting hesitation primarily as a reflection of integration and metaplanning.” (1988 : 377)

Which type of classroom exercise that would best promote fluency is not predicted from this. Some of the processing theories discussed, such as instance theory and the connectionist model, and chunking suggest that the mastery of formulaic patterns might promote fluency. This would suggest that conscious learning can lead to acquisition.

At the same time, the combination of controlled processing with automatic processing suggests that awareness of the language system is an important component in fluency. In addition to a communicative approach to language learning that might facilitate automatic processing, a strong language awareness component is equally necessary to foster the learner’s ability to make the best higher-level decision, to know among other things which issues of form that he or she can focus on. Assuming that grammatical proficiency includes the domains of pragmatics and discourse, this would include the speaker consciously focusing on appropriacy. A very basic example, would be deciding that a deleted relative pronoun with preposition stranding might be preferable in a conversation to a preposed prepositional wh-phrase (so-called ‘pied-piping’). That is not to say that this has to be a conscious decision. Presumably this will be part of the language acquirer’s competence. But surely it should be part of the speaker’s arsenal of information to be able to make informed judgements like this. Judging from my experience at universities in southern California, teachers still mark off for clause-final prepositions in compositions. And work by Biber has demonstrated that in formal texts it is very rare to find this.^{5} But perhaps this is now referring to a grammatically proficient rather than a fluent speaker.

Even this is still just half of the issue. In addition to language improvement, grammatical proficiency is important in a teacher’s repertoire of skills. In order for a teacher to appropriately edit student discourse, the teacher must continuously evolve in his/her understanding of the language system. For all of these reasons, the growth of grammatical proficiency is an integral component in professional development. Yet, this cannot be achieved by reading Swan from cover to cover. We have to remember two crucial points here as well: (i) grammar references contradict each other and provide false (due to simplification) or invalid (due to changes, dialects, conflicting standards, etc.) statements; (ii) as my colleague, Angi Malderez points out, we have only begun to describe the complexity of grammar in terms of rules and appropriacy (1993). In this connection, one might mention the point made by the editor of the Collins Cobuild dictionary and grammar, John Sinclair, who cites the example in the preface to the grammar of how sentences from a corpus, unlike made-up examples, might reveal new insights into the use of words, for example, that it is “bad things that break-out.”

Through the deliberations so far, we have been able to remind ourselves that grammar is a hypothesis, a tool for understanding and explaining. What is now required is to look at this understanding of grammar from the standpoint of language processing and then discuss the consequences for both theory and classroom.

3 Language Processing: Symbolic Knowledge, Linguistic Control and Pedagogy

At this point, I am going to recapitulate Bialystok’s findings in her 1990 publication Communication Strategies and relate them to the importance of grammatical proficiency and its communication in the classroom. A further tenet in her model is the similarity between L1 and L2. I will therefore also address a progression in levels of proficiency for L2 learners offered by Rutherford (1987), the use of coordination vs. subordination, and show how the parallel with native speaker progression helps to recognize an error as developmental. An issue that arises there is that of literacy and cognitive academic language aptitude [CALP, cf. Crawford 1989]. In the final part of the paper I will look at other samples of student error and show how they can be used in the classroom to heighten grammatical proficiency, and presumably thereby also enhance language proficiency and pedagogical awareness. The errors themselves first make sense from the type of overall framework proposed here.

3.1 The Bialystok Model

To start with Bialystok’s conclusions, the best way to increased proficiency is through advancing explicit awareness of the language system to enhance ‘symbolic knowledge’ and through practice to enhance ability to exercise control of linguistic processing. A growth in awareness means an increased flexibility, surely a key component in fluency, and greater sensitivity to the needs of the audience.^{6} For Bialystok these are skills reflective of a general maturation process. Certainly college educators sense the importance of the growth in these areas and many a corridor discussion takes place in searching for ways to foster them. Bialystok relates these as well to the increased ability to decontextualize knowledge, which she posits as advancing from implicit to explicit knowledge in the case of mastering a language.

Already a fact known to American ESL educators in thus appropriately acknowledged, namely the importance of literacy skills in the first language, or cognitive academic language proficiency ([CALP] Cummins) in mastering English as a second language in a school environment. Bialystok says for example that “learning to read […] requires that some of [the] knowledge of language has been made explicit and represented as symbolic knowledge” (1990 : 121). In discussing Cummins’ notion of CALP, Crawford observes that this is required by children “if they are to succeed in the context-reduced, cognitively demanding activities of reading, writing, mathematics, science and other school subjects” (Crawford 1989 : 107).

For Bialystok, the ability to objectify the language system, i.e. turn it into ‘symbolic knowledge’ is one of two basic strategies required to become proficient in a language. This is not necessarily a conscious activity. She draws a strong parallel here between L1 and L2 acquisition and indeed refuses to draw a distinction between the two. The development in this skill can take place via “self-reflection on knowledge” or “literacy instruction” (1990 : 124). There is also a give and take between this skill and that of linguistic control, which is a conscious focusing on form or “relevant and appropriate information” (op.cit. : 125).^{7} This latter skill is fostered by schooling as well as bilingualism. Bialystok cites numerous experimental work throughout to support her arguments. For example, one study showed that young people who were fluent but did not have schooling were not able to understand a syllogism (op.cit. : 128).

The conclusion to be drawn from the model is that making students aware of the language system, i.e. converting their knowledge to explicit knowledge, has a crucially important role to play in any academic language program. It is important from the point of view of increased language proficiency and from the pedagogical point of view. For both, “judging and correcting sentence grammaticality” depends “on the level of analysis of linguistic knowledge” (op.cit. : 122). This point would probably not need to be made if such emphasis had not been placed on acquisition vs. learning in recent years. From the Bialystok model, learning has a very important role to play in language proficiency. And it does not need to undermine the relevance of acquisition. For indeed in order to progress from implicit to explicit knowledge there has to be implicit knowledge!

Having established that focusing on rules of the language system can promote proficiency, the next issue is how this kind of explicit awareness can be achieved. Bialystok observes that the rules should be comprehensible “as an organizing principle of linguistic knowledge” (op.cit. : 125). Returning to an earlier point, it is not necessarily helpful to teach a reference grammar from cover to cover. In fact, this approach can have a negative impact if the learner gets lost in a sea of detail. Once again, the lessons of the evolution of generative grammar are appropriate here: principles are important that have a type of explanatory adequacy. It is also useful to recall at this point that linguistic theory is itself an amorphous mass seeking to make sense of the language system. Ideally, theory and practice can be merged so that each has an impact on the other. And this should take place at the teachertraining level.

Bialystok herself draws consequences from her study for communication strategies. It appears that the strategies themselves are not important, nor can they explicitly be taught. Rather, enhancing the two processing skills cited above will lead to increased flexibility on the student’s part so that the student will be able to adopt the right strategy from a number of choices in any given situation. Bialystok also cites Corder’s seminal distinction of message adjustment vs. resource expansion strategies. For her, most of the communication strategies have to do with ‘message adjustment’. Though she says that she agrees with Corder’s point of view that a balance should be achieved between the two, the model she proposes actually stresses the importance of message adjustment. In EFL literature, emphasis has been put on resource expansion as a risk-taking strategy that could advance competence (cf., for example, Medgyes 1989). Bialystok, however, argues that it is rather the challenge of the task that advances a learner’s competence. As this conclusion is based on her language-processing model, it is worth describing in greater detail how she arrives at this conclusion.

In discussing Váradi’s work she states that:

“The pivotal point of this system, then, is message adjustment. Once a speaker decides that the optimal message cannot be expressed, the speaker must then decide whether that meaning will be reduced or replaced by an adjusted meaning; when a meaning, either optimal or adjusted, cannot be directly expressed, the speaker must then decide whether expression of that message will be based on reduction or replacement of the usual forms.” (1990 : 33)

Though Bialystok puts the concept of message adjustment in a positive light, this is not the case with Váradi who says that it has “disturbing relevance for the classroom situation” (Váradi 1983 : 83). The replacement strategies for Váradi include circumlocution and paraphrase. In Bialystok’s system paraphrase includes approximation, word coinage and circumlocution, the three of which alone account for 93% of the utterances in her corpus from her study on “a group of 9-year old English speaking children learning French in immersion programmes.” (1990 : 58)

Meanwhile, these paraphrasing strategies are found among the resource expansion strategies in the systems of Faerch & Kasper and of Corder. These are the good strategies that teachers should foster while the message adjustment strategies are in Corder’s system ‘risk-avoidance’ strategies.

It becomes clear, then, that rather than stressing ‘achievement strategies’ (which have been equated with manipulation of expression), Bialystok is focusing on message adjustment strategies, which have been associated with ‘reduction strategies’ (and ‘replacement’ as well in Váradi’s system). The insight she offers is that the communication strategies studied have to do with manipulation of concepts, finding the attributes, etc. to communicate the concept even when the definitional term is not at hand. In her system, manipulation of form is a control-based strategy and accomplished predominantly through language transfer (op.cit. : 133), which takes up a very low percentage of the total strategies employed. In this situation, for her, the intention is held constant while the means of reference is altered (ibidem).

She quotes Kellerman as saying that “learners can either manipulate the concept so that it becomes expressible through available linguistic (or mimetic) resources, or they can manipulate the language so as to come as close as possible to expressing their original intention” (op.cit. : 111). The linguistic strategy involves “conscious transfer, foreignization, transliteration [… and] some instances of word coinage” (ibidem). Once again referring to Kellerman she states:

“the conceptual strategy invites the listener to infer the intended concept, providing hints about characteristics, category membership and the like. The code strategy points more directly at the object, restricting information to features of the object label.” (op.cit. : 115)

She acknowledges that in an experimental situation the participants are more reluctant to use coding strategies but presumably would not be in a normal conversation. This, then, seems to turn the old risk avoidance vs. achievement strategies upside down! If a language learner adopted a word, let’s say from the common L1 in a L2 classroom to express his thought, then he would very much avoid taking a risk.

For Bialystok, what really increases a learner’s achievement is the “extension and adaptation of resources to tasks that formally surpass a learner’s competence […]” (my emphasis) (op.cit. : 117). It is worth quoting at some length the rationale behind the emphasis on message-adjustment:

“The analysis based strategy is an attempt to convey the structure of the intended concept by making explicit the relational defining features. Speakers examine (not necessarily consciously) their symbolic representations of conceptual and linguistic structures in order to select features that will most accurately define the intended meaning. The strategies from the descriptive taxonomies that are included in the analysis-based strategy are circumlocution, paraphrase, transliteration and word coinage where the attempt is to incorporate distinctive features into the expression, and mime where the attempt is to convey important properties.” (op.cit. : 133)

Whether the language-processing model as described in Bialystok’s book will withstand the passage of time and the gathering of further evidence is not as critical as the sense of proportion that is offered through the model for the foreign-language classroom. By stressing both awareness of the language system, a type of monitoring and the challenge of the task as being crucially important in the language-learning process, Bialystok has offered the foreign-language instructor a ballast to the notions of acquisition and communicative tasks for advancing resources. Perhaps a way of describing this is to say that a foreign language instructor should challenge the student’s intellect, as expected from general pedagogy. This would seem to follow from the postulation that similar cognitive properties are involved in language learning as in other learning exercises.

3.2 Language Awareness

The preceding discussion lays the groundwork for analysing student sentences from a theoretical perspective that sheds light on the interlanguage revealed. The discussion will be restricted here to three topics: coordination vs. subordination, faulty argument structure and lexical subcategorization properties.

3.2.1 Coordination vs. Subordination

A pervasive problem that English students at ELTE have is the use of runon sentences (i.e. two or more clauses run together with no conjunction, punctuation or just a comma). Though these sentences seem to be a classic instance of interference, as these constructions are largely licit in written Hungarian, a closer look reveals that they can be viewed as developmental.

In the fourth chapter of Second Language Grammar, Rutherford points out the progression that second language learners make from the use of coordination to the use of subordination. In this progression he observes that the mistake of the cooccurrence of both the coordinator but and the subordinator although might be an example of a healthy error that shows the learner is making the transition to using subordination. It is readily apparent that an area has been touched on here that is central in composition courses where traditionally students have been trained to progress from parataxis to hypotaxis.

In reviewing 5–8-page papers from two university courses, a firstyear and a third-year course, with a total of about 27 students, I had to be very selective to limit my collection of run-on sentences to 40. Many of the sentences lend themselves readily to rewriting with subordination. Some of the most obvious examples are the following which are easily transformed into sentences containing participials, appositive NPs or relative clauses:

(5) He is always full of imagination, usually approaches tasks from different angles and he comes up with new ideas. [similar to a run-on]

(6) On the other hand, the boys are not interested in learning any languages at all, they consider it a waste of time.

(7) During most of the classroom time it is possible to make a healthy compromise, students can be given a good amount of time to produce things not for the sake of accuracy.

(8) Everybody is a member of a certain community, he possesses particular features of personality, he has his own experience and motivation, etc.

(9) By no means do all learners show interference errors in their speech, the amount always depends on the certain person.

(10) The mother tongue might have a facilitating effect where the systems of L1 and L2 resemble each other, it is the so-called positive transfer.

(11) The third group has no clear idea of going on, they do not know what their teacher wants to teach in the following month as obviously everything depends on her.

(12) Fathman wrote about the position of science and language learning. She points out that science and language learning are natural partners in the development of second language. [note the same problem here without the sentence being a run-on] (13) This is just one example, from the field of politics, it can be trans-

ferred to art, sports or whatever is happening in the world.

(14) Many people find foreign countries strange, they only feel comfortable in their own surroundings.

(15) They have an information processing system a mainly visually based one which is similar form as our verbal language, it definitely serves then some biological function.

(16) Jane Goodall, the famous scientist, spent more than thirty years observing chimpanzees in the wild in Tanzania (Gomba) she got very close to them, she made lifelong friends among them.

(17) One does not learn how to communicate nonverbal, there is no such subject in the school.

(18) On the other hand, females are said to be gossipy, they talk too freely and much in private situations.

(19) The advantage is that Esperanto is a neutral language, it does not have the nonverbal political and cultural base that all other natural languages have.

The same types of sentences might be set up in a composition course for native speakers as a sentence combining exercise. A factor in the grading of compositions, for example for the entrance exams for the English Department at the University of Budapest, is syntactic complexity. If a student has mastered the use of subordination and participials, his/her English is judged to be more ‘fluent’.

These sentences are often overlooked by my colleagues in the English Department. There is a feeling that this is an interference error and it is largely tolerated. (Students point out that elsewhere these sentences are not corrected in their compositions.) But the nature of the sentence construction changes if we view this as a developmental error, at least in equal proportions to being one of interference. (Note that this could also be overuse of subject in English. Hungarian is a pro-drop language.) And such a view suggests that their correction could play an important role in language awareness or in allowing the students to concentrate on form. Using the model of language processing outlined above, the student would then have at his/her disposal a greater degree of flexibility in choosing a syntactic structure for his/her thoughts. That is to say that the student would be able to exercise greater linguistic control, an important factor constituting fluency, in being able to focus on a choice of syntactic constructions.^{8} 3.2.2 Argument Structure

Hungarian and Portuguese university students I have worked with have a great problem with argument structure. The types of errors that they make are readily explicable from the generative framework. One example would be that of ergativity. Ergative verbs are clearly a marked class in English. The student tendency to use either the passive voice, subject–verb inversion or an expletive follows from Burzio’s generative model that posits the ergative subject in object position in the underlying structure.^{9} Without even considering the syntactic arguments put forward for this position, we can see that the model captures the intuition that the ergative subject is ‘object’ like in its semantic role. Though this would also be accounted for by saying that the thematic role of the subject is non-agentive, a patient or theme, the syntactic implications, for example for a second-language learner, are not as evident. I refer the reader to Burzio (1986), Rutherford (1987), Yip (1989), Zobl (1990) for a more detailed account of ergativity and its consequences for the second-language learner.

Rutherford (1987) posits the concept of ‘syntactic-semantic’ distance to capture the type of difficulty second language learners of English have in confronting differences between underlying argument structure and surface word order. The following two sentences give evidence of how students grapple with the problem of so-called tough-movement:

(20) Summing up, he was a teacher, a scientist and a monk in one person, each of them is difficult to fulfill its requirements during a lifetime.

(21) However, as these types of relative clauses are commonly introduced by the pronoun which, it is irrelevant to be dealt with it here. On the other hand, nominal relative clauses are very important to be discovered.

In sentence (20) the student has marked the NP at both its surface subject position and underlying object position (the idea of the sentence is ‘each of these requirements is difficult to fulfill’). Rutherford notes that students progress from sentence (b) to sentence (c) below in their acquisition of the language:

(22) a. To imitate his accent is impossible.

b. It’s impossible to imitate his accent. c. His accent is impossible to imitate.

Sentence (22a) contains a sentential subject, a construction which has proven to be harder to process.^{10} Sentence (22b) has the argument structure intact, in that the object his accent follows the head to which it is a complement, imitate. In sentence (22c) the head and the complement are now separated. A phenomenon that we find in second-language data is the retention of a pronoun in the original object position, a sentence like:

(22) d. His accent is impossible to imitate it.

This could be another ‘healthy’ error as the student is working out the rule for tough-movement. At the same time a deeper look at the syntax makes the source of the mistake abundantly clear as a predictable developmental error. As a student of mine once put it, “the language invites the learner to make certain errors.” This might be a way of defining ‘markedness.’

So in sentence (21) above we might say that the student has really played it safe: we have a subject that could be interpreted either as an expletive or a content pronoun (assuming the acceptability of it is irrelevant to deal with here), the passive voice to show that the subject is thematic and related to an object position and then the original object position explicitly marked by the pronoun. Once again from a generative perspective and in terms of interlanguage all of this makes perfect sense. Without this tool it would appear that the student is engaging in utter chaos!

The following three sentences are revealing in showing the type of difficulty the student can steer into with argument structure. Sentences (23) and (24) contain verbs that can either be ergative or transitive with an agentive subject. The ergative use is intended by the student, but the passive voice is chosen for both. This is a typical error for Chinese students as pointed out by Yip (1989) but apparently not for Portuguese students, who produce the other two predicted constructions mentioned above (cf. Shanklin 1992, 1993). In (25), a verb that is not ergative and must be used with the passive voice when there is a thematic subject is treated like an ergative verb:

(23) If something is compulsory, it’ll never work the way it should be worked.

(24) I have chosen this statement to be my closing motto because this leads us back to the question of awareness and where my string of thoughts were started.

(25) The third word have implanted to Hungarian so well, that both its pronunciation and spelling have altered, however its strangeness is still very conspicuous.

These errors can be seen to be developmental. Even advanced ELTE students had great difficulty in identifying the error in sentence (25). At the same time, I wonder whether English is not becoming increasingly an ‘ergative’ language, in that it seems to me that more and more verbs that are typically transitive tend to be used intransitively. Could one say perhaps, for example, ‘the space shuttle launched, but there were problems in the electronic system’? I have to start paying attention to when a simple intransitive is used when I would otherwise expect the passive voice. This is for me the main point about language awareness. I came across such an example in a recent Newsweek article (“Tailhook’s ‘Lightning Rod’ ”, 28 Feb. 1994, p.35). There the sentence appears:

(26) She denies it all and the rumors have never checked out.

Note the difference between this ergative construction and the passive. In the passive equivalent no one looked to see whether the rumors were true, whereas in the sentence above, the rumors were looked into but not validated. Some of the students I asked got this meaning of the ergative but most did not. If they looked in Collins Cobuild, or the Collins Cobuild Dictionary of Phrasal Verbs, or in Webster’s New World they would find no mention of the ergative construction used here. The usage is mentioned in Longman’s Dictionary of Phrasal Verbs.

3.2.3 Subcategorization Properties^{11}

It also seems to me that a verb can change its subcategorization properties. I have spent years telling students that you cannot say I explained him the answer the way would can I told him the answer. I am becoming less sure and wonder whether eventually the subcategorization property of explain will not become identical to tell. I thought that agree could not be used transitively except as a prepositional verb, while on BBC the construction they agreed the plan is quite standard.^{12} The theory provides for this, in that the subcategorization properties of lexical items are the critical point in deciding arbitrary features of single language usage. Perhaps one day all natives will join many language learners in saying I suggested them to go even if we might wrinkle up our noses currently.

But once again the only way out is a continuous on-going evaluation of the language. One of my students, a bilingual speaker of Hungarian and English thought that the construction in (27) would be acceptable:

(27) The crack was failed to be noticed.

This was an attempt to put a sentence from a worksheet he failed to notice the crack into the passive. I thought that this sentence was unacceptable and that the appropriate passive would be the crack failed to be noticed but looked for a means to explain my position. First I contrasted fail with try, pointing out that one can say the crops failed but not the crops tried, a way of showing that fail can be an ergative verb whereas try cannot. I then contrasted fail with expect and produced the following three sets of sentences:

(28) a. John tried to see him.

b. ?The crops tried to grow. c. *John was tried to see him. d. *The crops were tried grow.

(29) a. John failed to see him.

b. The crops failed to grow. c. *John was failed to see him. d. *The crops were failed to grow.

(30) a. John expected to see him.

b. *The crops expected to grow. c. John was expected to see him. d. The crops were expected to grow.

The difference between (28b) and (29/30b) is simply that expect and try cannot be used ergatively whereas the contrast between (30c/d) and (28/29c) and (28/29d) is due to the fact that expect can have two different subjects in the main clause and complement clause whereas fail and try cannot have two separate subjects (*He failed/tried John to pass the exam.). All of this is readily explainable but intricate! The real question I suppose comes back to how one’s knowledge of the language system can be taught, how a skill can be imparted. I would maintain that the attitude towards grammar advocated in this paper is above all an approach and that this approach can be communicated. I think all to often educators associate the term ‘generative grammar’ with an excessive, perhaps esoteric, formalism. But that doesn’t have to be the case. One could claim that theoretical linguistics are busy concerning themselves with the formalism while the results of the research are implementable in a far less complex manner.

Fortunately at about the same time that I was discussing this grammar point with a first-year near native speaker language improvement course, I came across the following sentence in a linguistics article:

(31) Thus obvious generalizations fail to be expressed. (Sag et al., 1984)

Finally, I would like to contrast two sentences from two first-year papers in an Introduction to Linguistics course:

(32) It is also very interesting to look at that human languages have more tokens for one meaning. [1/NG] (33) From my point of view, however, it is more important to take a look

at the fact how necessary these factors are. [1/NR]

These two sentences throw up interesting reflections about what constitutes the subcategorization properties of the noun and verb look at. In (32) we can say that a preposition does not take a finite clause for its complement. In a sense the insertion of the fact that is a syntactic device to get around this constraint. In (33) on the other hand we see a manifestation of the curious property that a preposition can take a wh-clause as its complement. Therefore the syntactic device the fact that has been introduced unnecessarily and inappropriately. The first sentence is from a native speaker of German and the second from a native speaker of Russian. Interference is not the key to what is going on, while the subcategorization properties of prepositions are. So once again we are focused on developmental errors and an awareness of the language system allows us to recognize and repair the errors, identify the source and hopefully, to end on an optimistic note, to communicate this information and this skill which can only assist in improving both language proficiency and pedagogical skills. For the method of communicating this skill, I would simply refer to the approach taken in this paper.

Finally, a question is left for theoretical linguists that I am sure is being worked on: just what are the deeper principles at work in the language that dictate that prepositions have the curious subcategorization properties discussed above. My colleague, Mark Newson has suggested that wh-clauses are more ‘nominal-like’ than that-clauses, which would account for their categorizing with NPs while that-clauses do not (p.c.). An investigation of this would involve an analysis of the Complementizer Phrase (CP) in X-bar syntax.

4 Conclusion

The purpose of the paper has been to provide a coherent framework for the understanding of grammar and its role in the second-language classroom. In the final sections, examples were given for how student error, from the vantage point of this conceptual framework, can be used to advance understanding of the language system. This understanding of the language system has been defined as grammatical proficiency. The notion of grammar is seen to be of far greater complexity than simply the review of a pedagogic grammar or the teaching of structures of progressive difficulty.

Crucial for this understanding has been the notion of interlanguage, a generative term that puts the grammar to be investigated in the mind. Putting it there also raises the issue of language processing skills.

Through the investigation, some basic notions of a communicative language curriculum have been reassessed: the fluency vs. accuracy distinction, the input hypothesis, the separation of acquisition vs. learning and stressing the expansion of linguistic resources over message adjustment. It was suggested that awareness of the language system, and therefore an aspect of language proficiency associated with monitoring or accuracy, could be a factor in fluency. Secondly, Bialystok’s work on the relationship between general cognitive properties and language skills suggests that Krashen’s model of the affective filter preventing the language acquisition device from working to full capacity in the adult learner is at best inadequate. Though motivation is an important factor in successful language learning, Bygate’s admonition to reflect on the impact of an activity on language processing skills is equally valid.^{13} Finally, there are grounds for arguing that confronting language learners with challenging tasks that force them into a deeper conceptual and linguistic analysis will contribute to language proficiency. By doing so, the important cognitive and communicative properties of flexibility and developing sensitivity to the listener’s needs are fostered. Indeed, there would be a similarity between what goes on in a language-learning classroom and the otherwise generally acknowledged need to push students and ourselves beyond our narrow, ethnocentrically constrained viewpoint to appreciate other points of view.

None of this is intended to disparage those concepts and distinctions, which are important in a teacher-training program. These notions need, however, to be reassessed and relativized in a more advanced languagelearning model. This is reminiscent of the point made by Burton-Roberts about the need for different models or analyses in different stages of the learning process. In fact, we might call the whole enterprise: Beyond Krashen. The importance of this for the language teacher is that in essence every language teacher has some model of language processing that s/he is operating with, implicit and explicit assumptions about how language acquisition occurs. The applied linguist has an important role to play in making these assumptions explicit. In a teacher training program an elucidation of potential models should be a key component in the curriculum. Finally the reflective teacher will elaborate and modify such models, feeding back into the theory behind them.

notes

* This paper was used as the basis for a workshop at the 28th International IATEFL Conference in Brighton, England, 6–9 April, 1994. For the working through of the ideas in this paper I am first of all indebted to my team of colleagues at the Centre for English Teacher Training (CETT) and the former Department of English Language and Literature (DELL) at Eötvös Loránd University (ELTE) in Budapest. I would also like to acknowledge my degree of indebtedness to William Rutherford at the University of Southern California. In addition, Enikő Csomay has been an invaluable source of ideas and material in the field of applied linguistics. I suppose my greatest debt is to my students, who always have the best ideas.

[1] For the EFL teacher, the interaction would conform to the model proposed by Wallace to depict the process of a reflective teacher, who brings in theoretical knowledge and teaching skills into the classroom and while solving practical classroom problems reflects on this ‘received and experiential knowledge’ “in the context of professional action (practice)” (1991 : 56). For the theoretical linguist the interaction would conform to the type that according to Goldstine (1972) has always driven science forward:

“It is worth recalling that prior to this time the state of mathematics in Europe was not substantially more advanced than that in the Arab world, based as it was on European and Chinese ideas and concepts. Then suddenly, as a result of a bringing together of mathematics and physics, something happened in Europe that started science on the path that led from Galileo to Newton. This melding of practical and empirical knowledge with mathematics was the magic touchstone.”

[2] According to Stern, language proficiency, “interpreted as communicative competence (including linguistic competence) […] means an ability to use the language without giving linguistic forms, rules and meanings any specific thought” (1992 : 72). Once again, I would say the skill component of linguistic competence would entail the conscious focusing on form.

[3] Burton-Roberts also makes reference to this in Analysing Sentences. In the

introduction to that work he notes:

“[…] I have concentrated on presenting a single, more or less traditional, analysis of each structure considered, without overburdening the reader with too much discussion of how that analysis might or might not be justified in light of further evidence. This might give the misleading impression that there is just one possible analysis and that there is universal agreement that it is the one in this book! This is far from being the case. But sometimes the evidence that might support an alternative analysis is complex and indirect and its discussion would be inappropriate in such an introduction. The reader should bear in mind, then, that we are never irrevocably committed to a particular analysis but are free to amend it in light of further evidence.” (1986 : 4)

In addition to the point about accessibility, this presupposes the view that grammar is understood as a hypothesis about the language and in a Popperian spirit, the task of a linguist is to prove his initial hypotheses wrong on the way to an increasingly sophisticated understanding of the language system. Professor László Varga at ELTE feels that the use of the verbal group is better for students as it helps them better access available English grammars that use this or a like concept.

[4] On BBC the construction was used none of their lives is in danger (6 Feb. 1994). My first year class investigated the construction in grammars and found the distinction between notional and grammatical concord in Quirk et al., no mention in Thomson & Martinet, both singular and plural use without distinction in Swan and the listing of none of as a quantifier in Collins Cobuild with the accompanying explanation that the verb is plural if a plural noun follows.

[5] Cf. Biber (1988) and a discussion about this in E. Finegan & N. Besnier (1989), Chapter 13, Registers.

[6] About flexibility, Bialystok observes at the end of her book: “The more language the learner knows, the more possibilities exist for the system to be flexible and to adjust itself to meet the demands of the learner” (1990 : 147). Earlier she cites Sternberg’s theory of cognitive development in which “development is characterized by ‘greater flexibility and more appropriate strategy or information utilization […] with age’ to reach those goals” (op.cit. : 9). Concerning the needs of the audience, Bialystok cites one study (Whitehurst & Sonnenschein 1983) on L1 acquisition that claims that “children […] begin by acquiring global communication rules. The rules and procedures they develop become increasingly specific, moving on, for example, to rules for maintaining sensitivity to listener” (op.cit. : 9). In another study, Shatz states that “true communication occurs when the sender takes the receiver’s capacity to understand into account and exploits multiple sources of information to communicate these properties to a listener” (1983 : 97).

[7] ‘Control of linguistic processing’ is defined as “the ability to control attention to relevant and appropriate information and to integrate those forms in real time” (Bialystok 1990 : 125).

[8] Note that the environment for the run-on sentence typically involves a subject pronoun that repeats the subject of the first clause.

[9] That is to say that a verb like happen would have the underlying structure happens the war from which the surface structure the war happens is derived. The error is much more noticeable if the ergative verb is uniquely intransitive, like happen, rather than having a transitive counterpart, like break. Thus a student might use the expression the cup has been broken and mean the cup broke, which has no implied agentive.

[10] Cf. L. Frazier & K. Rainer (1988).

[11] In the Principles and Parameters framework, much of the grammatical information has been put in the (mental) lexicon. Especially arbitrary features should be dealt with as idiosyncratic lexical features and separate from more fundamental principles. This is the spirit, for example, of Radford (1988).

[12] One could easily write a study on the BBC usage of English. Another construction that I frequently hear is in ten minutes from now. Concerning the subcategorization property of the verb agree, it is perhaps interesting to note that whereas no mention of the construction alluded to (agree as a monotransitive rather than prepositional verb) is made in the BBC Dictionary, it is mentioned in Collins Cobuild.

[13] For a critique of the input hypothesis along lines that would at least mesh with those here, cf. Bley-Vroman (1989).

references

Bialystok, E. (1990) Communication Strategies: A Psychological Analysis of Second-Language Use. Oxford: Basil Blackwell.

Biber, D. (1988) Variation across Speech and Writing. Cambridge: Cambridge University Press.

Bley-Vroman, R. (1989) ‘What is the logical problem of foreign language learning?’ In: Gass & Schachter 1989 : 41–68.

Burton-Roberts, N. (1986) Analysing Sentences. London: Longman.

Burton-Roberts, N. (1993) ‘Use, mention, and linguistic occurrence: a representational approach’. lecture held at Eötvös Loránd University, Budapest, 8–12 November 1993.

Burzio, L. (1986) Italian Syntax: A Government-Binding Approach. Dordrecht: D. Reidel Publishing Co.

Bygate, M. (1988) ‘Units of oral expression in language learning in small group interaction’. In: Applied Linguistics 9:1. 59–82.

Canale, M. (1983) ‘From communicative competence to communicative language pedagogy’. In: J. C. Richards and R. W. Schmidt, eds. Language and Communication. London: Longman. 2–27.

Chaudron, C. (1988) Second Language Classrooms: Research on Teaching and Learning. Cambridge: CUP.

Corder, S.P. (1983) ‘Strategies of communication’. In: C. Faerch and G. Kasper 1983 : 15– 19.

Chomsky, N. (1986) Knowledge of Language. New York: Praeger.

Crawford, J.(1989) Bilingual Education: History, Politics, Theory and Practice. Trenton, NJ: Crane Publishing Co.

Faerch, C. and G. Kasper (1983) ‘Plans and strategies in foreign language communication’. In: C. Faerch and G. Kasper 1983 : 20–60.

Faerch, C. and G. Kasper, eds. (1983) Strategies in Interlanguage Communication. London: Longman.

Finegan, E. and N. Besnier (1989) Language its Structure and Use. San Diego, CA: Harcourt, Brace and Jovanovich.

Frazier, L. and K. Rainer (1988) ‘Parameterizing the language processing system: leftvs. right-branching within and across languages’. In: J. Hawkins, ed. Explaining Language Universals. Oxford: Blackwell. 247–279.

Gass, S. M. and J. Schachter, eds. (1989) Linguistic Perspectives on Second Language Acquisition. Cambridge: CUP.

Goldstine, H. (1972) The Computer from Pascal to Neumann. Princeton, NJ: Princeton University Press.

Greenbaum, S. and R. Quirk (1990) A Student’s Grammar of the English Language. London: Longman.

James, C. (1993) ‘What is applied linguistics’. In: International Journal of Applied Linguistics 3:1. 17–32.

Malderez, A. (1993) ‘Communicative Competence’. Applied Linguistics Lecture Series. Fall 1993, CETT.

Medgyes, P. (1989) ‘The correction of covert and overt error’. In: Annales Universitatis Scientiarum Budapestinensis de Rolando Eötvös Nominatae, Sectio Linguistica, Tomus XX. 179–193.

Medgyes, P. (1994) The Non-native Teacher. Houndmills and London: Macmillan Co.

Radford, A. (1988) Transformational Grammar. Cambridge: CUP.

Rutherford, W. E. (1987) Second Language Grammar: Learning and Teaching. London and New York: Longman.

Quirk, R., S. Greenbaum, G. Leech and J. Svartvik (1985) A Comprehensive Grammar of the English Language. London and New York: Longman.

Sag, I., G. Gazdar, T. Wasow and S. Weisler (1984) ‘Coordination and how to distinguish categories’. In: Natural Language and Linguistic Theory 3. 117–171.

Schmidt, R. W. (1992) ‘Psychological mechanisms underlying second language fluency’. In: Studies in Second Language Acquisition 14. 357–385.

174 M. Trevor Shanklin

Shanklin, M. T. (1992) ‘Theseus out of the Labyrinth’. Unpublished ms., presented at 22nd Linguistic Symposium on Romance Linguistics, University of Texas at El Paso and Universidad Autonoma de Ciudad Juarez, 20–23 Feb. 1992.

Shanklin, M. T. (1993) ‘Student interlanguage from a syntactic point of view’. Unpublished ms., presented to the Theoretical Linguistics Department of the Eötvös Loránd University, Budapest, 26 Feb. 1993.

Stern, H. H. (1992) Issues and Options in Language Teaching. Oxford: OUP. Swan, M. (1980) Practical English Usage. Oxford: OUP. Taylor, D. (1988) ‘The meaning and use of the term “competence” in linguistics and

applied linguistics’. In: Applied Linguistics 9:2. 148–168. Thomson, A. J. and A. V. Martinet (1986) A Practical English Grammar, 4th ed. Oxford:

OUP. Turóczi, A. (1994) ‘Gift for languages’. Unpublished B.ed. thesis, Centre for English

Teacher Training, Eötvös Loránd University, Budapest. Yip, V. (1989) ‘Aspects of Chinese/English interlanguage: syntax, semantics and learn-

ability’. Unpublished doctoral dissertation, University of Southern California, Los Angeles, CA.

Váradi, T. (1983) ‘Strategies of target language learner communication: message adjust-

ment’. In: C. Faerch and G. Kasper 1983 : 79–99. Wallace, M.J. (1991) Training Foreign Language Teachers. Cambridge: CUP. Zobl, H. (1989) ‘Canonical typological structures and ergativity’. In: Gass & Schachter

1989 : 203–221.

— (1992) BBC English Dictionary. London: BBC English and Harper Collins.

— (1987) Collins Cobuild English Language Dictionary London: Harper Collins.

— (1990) Collins Cobuild English Grammar London: Harper Collins.

— (1989) Collins Cobuild Dictionary of Phrasal Verbs. London: William Collins Sons and Co. Ltd.

— (1983) Longman Dictionary of Phrasal Verbs. London: Longman.

— (1988) Webster’s New World Dictionary, Third College Edition. Cleveland, Ohio: Simon and Schuster, Inc.